11) You have a 55 year old male in LTC facility that has a PEG tube. Height is 6 land weight is 175 lbs. The patient is getting enteral feeding. You perform all the necessary calculations and determine his BEE is 1755 Kcal. You do an adjustment to the BEE of 1.2. You do an adjustment to the Protein requirements of 1.3. You decide to use Boost Plus which is 1.52 Kcal /ml. It has 78% free water in each 237ml can. 16% of the calories are protein according to the label. How much Additional water do you have to add to meet the patients daily fluid needs for this patient??
Answer #1

Daily basic energy expenditure of patient = 1755kcal

Thermal effect of food = 1.2 (which was adjusted by you)

Resultant energy requirement = BEE * 1.2=1755 * 1.2=2106kcal

Total free water in boost plus is 78%

Amount of water in ml = (237 *78)/100= 184.9= 185 ml of water is present in a 237 ml of boost plus.

Additional water requirements can be calculated by multiplying the patient weight in kilograms by 35 ml/kg= 79kg * 35ml/kg=2765 ml of fluid

So the extra water needed is 2765-2106 =656 ml water he needs additional .